# Excitation function of femtoscopic Lévy source parameters of pion pairs in EPOS4

### v1

@@ Abstract
Three-dimensional (3D) femtoscopic source parameters of pions provide a sensitive probe of the space-time structure of particle-emitting sources in high-energy heavy-ion collisions. Compared to one-dimensional measurements, 3D femtoscopy not only provides a valuable cross-check but also offers a more complete characterization of the source geometry and its dynamical evolution. Particularly, differences between the "out" and "side" directions are sensitive to signals of a strong first-order phase transition, while the collision-energy dependence of Lévy radii may reveal non-monotonic features related to the equation of state (EoS).
In this work, we systematically investigate the transverse mass and collision-energy dependence of the three-dimensional femtoscopic parameters of pion pairs with Lévy-type sources in the STAR Beam Energy Scan (BES) range from $\sqrt{s_{NN}}=7.7$ to $200$ GeV using the EPOS4 model. The analyzed parameters include the Lévy index $\alpha$, the correlation strength $\lambda$, and the three-dimensional radii $R_{\mathrm{out}}$, $R_{\mathrm{side}}$ and $R_{\mathrm{long}}$ , derived quantities such as the radius difference $R_{\mathrm{diff}}=R_{\mathrm{out}}^{2}-R_{\mathrm{side}}^{2}$ and the ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$ are also investigated. The results show that the extracted radii $R_{\mathrm{side}}$ and $R_{\mathrm{long}}$ decrease with increasing transverse momentum and increase gradually with collision energy, while $R_{\mathrm{out}}$ shows little energy dependence. The Lévy index $\alpha$ exhibits only a mild dependence on $m_{T}$ and collision energy, whereas the correlation strength $\lambda$ shows a clear $m_{T}$ dependence and generally decreases with increasing collision energy. A comparison with EPOS3 results indicates general agreement within approximately $2\sigma$, with the notable exception of $R_{\mathrm{side}}$, which is systematically smaller in EPOS4.
*Corresponding author(s). E-mail(s):huang@ttk.elte.hu
@@ 1 Introduction
Traditionally, the source distribution was modeled by a Gaussian form. However, in recent years, various experiments–from SPS [1, 2, 3], through RHIC [4, 5, 6, 7, 8], to LHC [9, 10]–have shown that Lévy-stable source distributions provide a better description of the measured correlation functions, particularly in cases where anomalous diffusion or long tails appear in the emission profile. Several physics mechanisms in high-energy heavy-ion collisions may contribute to such non-Gaussian behavior, including critical phenomena, jet fragmentation, and anomalous diffusion [11, 12, 13, 14]. The ELTE group has recently developed and applied a three-dimensional (3D) Lévy femtoscopy method to investigate pion sources in relativistic heavy-ion collisions [15]. This study provided a comprehensive framework for extracting the Lévy index of stability ($\alpha$), source radii ($R_{\rm{out}},R_{\rm{side}},R_{\rm{long}}$), and correlation strength ($\lambda$), offering improved sensitivity to non-Gaussian features of the source.
A particularly intriguing aspect of Lévy femtoscopy is its potential to reveal signatures of the QCD critical point. Near the critical point, long-range correlations are expected, leading to a non-trivial, possibly non-monotonic energy dependence of the Lévy index $\alpha$ [16]. Motivated by this, we extend our previous work by systematically analyzing the energy dependence of femtoscopic source parameters within the EPOS4 event generator framework. This provides a theoretical baseline for interpreting future experimental data and for exploring possible critical phenomena.
@@ 2.1 Distance distribution measurement and extraction of key parameters
In simulations, the pair distance distribution $D(\vec{\rho})$ can be directly obtained, where $\vec{\rho}$ represents the three-dimensional spatial separation vector of the particle pairs. This allows the extraction of key source parameters such as the Lévy index $\alpha$, the source size $R$, and the correlation strength $\lambda$. In experimental measurements, however, the spatial distributions are not directly accessible. Instead, one measures the correlation function in momentum space (as a function of momentum difference $\vec{q}$), which encodes information about both the spatial distribution $D(\vec{\rho})$ and the pair wave function $\psi_{\vec{q}}(\vec{\rho})$:
@@
where $\vec{\zeta}$ is the integration variable, the superscript $T$ denotes the transpose, while $R^{2}=\mathrm{diag}\left(R_{\mathrm{out}}^{2},R_{\mathrm{side}}^{2},R_{\mathrm{long}}^{2}\right)$ denotes the diagonal matrix of Lévy-scale parameters defined in the Bertsch–Pratt “out-side-long” coordinate system [20, 21], neglecting cross-terms. The $2^{2/\alpha}$ factor appears due to $D$ being defined as the autoconvolution of the Lévy-distributed single-particle source.
In this study, for each energy, a total of 4000 minimum-bias events were generated using EPOS4 with its default X3F Equation of State with cross-over transition and three flavour conservation [22, 23]. Unlike Ref. [24], where the angle-averaged distance distribution $D(\rho)=\frac{1}{4\pi}\int D(\vec{\rho})\,d\Omega$ was measured, but similarly to Refs. [25, 15, 26], we investigate the 3D pair distribution. To express the source coordinates in the Bertsch-Pratt coordinates of the longitudinally co-moving system (LCMS) [27, 28, 29], the components of $\vec{\rho}$ are computed from the particle positions in the laboratory frame as follows:
@@
Here, the azimuthal angle $\varphi$ is defined by $\cos\varphi={K_{x}}/{k_{T}}$, $K_{0}$ and $(K_{x},K_{y},K_{z})$ denote the temporal and spatial components of the pair’s average four-momentum $\vec{K}$, $k_{T}=\sqrt{K_{x}^{2}+K_{y}^{2}}$ is the pair transverse momentum, and $(r_{x},r_{y},r_{z},t)$ are the particle coordinates in the laboratory frame. Detailed derivations and conventions can be found in the Methods sections of Refs. [15, 25].
For each event, the distribution of relative distances $D(\vec{\rho})$ between pion pairs was constructed in the longitudinally co-moving system. The distance distributions were then fitted using the Lévy parameterization, with the fitting procedure following the methodology outlined in Ref. [15], to extract the parameters $\alpha$, $R_{\rm{out}}$, $R_{\rm{side}}$, $R_{\rm{long}}$, and $\lambda$. The fits were found to describe the obtained sources well in all transverse mass bins at all collision energies. Thus, the interpretation of the extracted Lévy source parameters is adequate. As discussed in Ref. [25], the fitted distribution corresponds to the core-core source, consisting of pairs of pions produced close to the interaction point, i.e., primordial pions and the decay products of the shortest lived resonances.
@@ 2.2 Systematic uncertainties
The systematic uncertainties were estimated by varying several analysis conditions (sets) around the default configuration. The corresponding cuts applied for different sources are summarized below, with the numbers in parentheses indicating the values used for the various collision energies.
(1) Number of averaged events
Due to limited statistics in single-event analyses, events were merged to achieve sufficient pair statistics. It was carefully investigated and confirmed that the extracted parameters converge at a sufficiently large event number. For different collision energies, the number of merged events ($N_{\rm events}$) was varied as follows (the middle value is taken as default):
@@
  • 130 and 200 GeV: (70, 100, 130)
(2) Pair momentum difference
In quantum-statistical correlations, the signal appears at limited values of relative momentum $Q_{\mathrm{LCMS}}$. In order to conceptually replicate experimental procedures, we utilize pairs up to a maximal momentum difference of $Q_{\mathrm{LCMS}}^{\mathrm{max}}=c\cdot\sqrt{m_{T}}({\rm MeV})$ where the coefficient $c$ was varied for different collision energies (again the middle value is the default):
@@
(3) Fit range in $\rho$
The stability of the extracted source parameters was examined by varying the fitting range in $\rho$:
@@
(4) Centrality definition
Centrality was defined using either the impact parameter (bimbim) from the EPOS4 definition or the charged-particle multiplicity ($N_{\mathrm{ch}}$). The two-dimensional ($bim,N_{\mathrm{ch}}$) distributions of the dataset were analyzed to define centrality bins according to the accumulated event fractions (0-10%, 10-20%, etc.). The centrality boundaries extracted from the impact parameter distribution are further found to be consistent with those provided directly by EPOS4. Overall, the choice of centrality definition contributes less than 2% to the systematic uncertainty of all extracted femtoscopic quantities.
@@
$\sqrt{s_{{}_{NN}}}$ Source $\alpha$ $\lambda$ $R_{out}$ $R_{side}$ $R_{long}$
7.7 GeV $N_{\rm events}$ $0.59\%$ $0.25\%$ $0.24\%$ $0.28\%$ $0.45\%$
$Q_{\mathrm{LCMS}}^{\rm max}$ $0.66\%$ $0.13\%$ $0.17\%$ $0.98\%$ $2.15\%$
$\rho_{\rm fit}^{\rm min/max}$ $0.38\%$ $0.16\%$ $1.01\%$ $0.12\%$ $0.69\%$
Centrality $0.89\%$ $0.23\%$ $0.14\%$ $0.58\%$ $0.76\%$
Total $1.31\%$ $0.39\%$ $06\% 1.1$$18\%$ $2.43\%$
19.6 GeV $N_{\rm events}$ $0.30\%$ $14\% 0.0$$1\% 0.0$$09\%$ $0.16\%$
$Q_{\mathrm{LCMS}}^{\rm max}$ $48\% 0.0$$03\%$ $0.43\%$ $1.02\%$ $1.73\%$
$\rho_{\rm fit}^{\rm min/max}$ $1.34\%$ $0.36\%$ $43\% 0.0$$08\%$ $0.79\%$
Centrality $0.51\%$ $0.24\%$ $0.41\%$ $0.51\%$ $0.23\%$
Total $1.54\%$ $0.45\%$ $0.73\%$ $1.14\%$ $1.92\%$
200.0 GeV $N_{\rm events}$ $11\% 0.0$$06\%$ $13\% 0.0$$7\% 0.0$$06\%$
$Q_{\mathrm{LCMS}}^{\rm max}$ $2\% 0.440$$4\%$ $0.24\%$ $1.03\%$ $1.95\%$
$\rho_{\rm fit}^{\rm min/max}$ $0.53\%$ $0.14\%$ $0.93\%$ $0.27\%$ $0.85\%$
Centrality $\% 0.170.17\%$ $0.59\%$ $0.40\%$ $0.34\%$
Total $0.70\%$ $0.50\%$ $\% 1.141.14\%$ $2.16\%$
@@ 3 Results and discussions
In this section we present the results for the femtoscopic source parameters as functions of collision energy, focusing on three main aspects: source shape ($\alpha$), source size ($R_{o,s,l}$), and correlation strength ($\lambda$).

### v2

@@ Abstract
Three-dimensional (3D) femtoscopic source parameters of pions provide a sensitive probe of the space-time structure of particle-emitting sources in high-energy heavy-ion collisions. Compared to one-dimensional measurements, three-dimensional femtoscopy not only provides a valuable cross-check but also offers a more complete characterization of the source geometry and its dynamical evolution. Particularly, differences between the “out” and “side” directions are sensitive to signals of a strong first-order phase transition, while the collision-energy dependence of Lévy radii may reveal non-monotonic features related to the equation of state.
In this work, we systematically investigate the transverse mass ($m_{T}$) and collision-energy ($\sqrt{s_{NN}}$ ) dependence of the three-dimensional femtoscopic parameters of pion pairs with Lévy-type sources in the STAR Beam Energy Scan (BES) range from $\sqrt{s_{NN}}=7.7$ to $200$ GeV using the EPOS4 model. The analyzed parameters include the Lévy index $\alpha$, the correlation strength $\lambda$, and the three-dimensional radii $R_{\mathrm{out}}$, $R_{\mathrm{side}}$ and $R_{\mathrm{long}}$, corresponding to the outward, sideward, and longitudinal (beam) directions. Derived quantities such as the radius difference $R_{\mathrm{diff}}=R_{\mathrm{out}}^{2}-R_{\mathrm{side}}^{2}$ and the ratio $R_{\mathrm{out}}/R_{\mathrm{side}}$ are also investigated. The results show that the extracted radii $R_{\mathrm{side}}$ and $R_{\mathrm{long}}$ decrease with increasing transverse mass and increase gradually with collision energy, while $R_{\mathrm{out}}$ shows little energy dependence. The Lévy index $\alpha$ exhibits only a mild dependence on $m_{T}$ and collision energy, whereas the correlation strength $\lambda$ shows a clear $m_{T}$ dependence and generally decreases with increasing collision energy. A comparison with EPOS3 results indicates general agreement within approximately $2\sigma$, with the notable exception of $R_{\mathrm{side}}$, which is systematically smaller in EPOS4.
keywords:
HBT femtoscopy , Lévy source , EPOS4 , Excitation function
journal: Physics Letters B
[…]
[ELTE]organization=Department of Atomic Physics, ELTE Eötvös Loránd University,addressline=Pázmány Péter sétány 1/A, city=Budapest, postcode=H-1117, state=, country=Hungary
@@ 1 Introduction
Traditionally, the source distribution was modeled by a Gaussian form. However, in recent years, various experiments–from SPS [1, 2, 3], through RHIC [4, 5, 6, 7, 8], to LHC [9, 10]–have shown that Lévy-stable source distributions provide a better description of the measured correlation functions, particularly in cases where anomalous diffusion or long tails appear in the emission profile. Several physics mechanisms in high-energy heavy-ion collisions may contribute to such non-Gaussian behavior, including critical phenomena, jet fragmentation, and anomalous diffusion associated with the long-range transport in the expanding medium [11, 12, 13, 14]. A three-dimensional (3D) Lévy femtoscopy method has been recently developed and applied to investigate pion sources in relativistic heavy-ion collisions [15]. This study provided a comprehensive framework for extracting the Lévy index of stability, source radii, and correlation strength, offering improved sensitivity to non-Gaussian features of the source.
A particularly intriguing aspect of Lévy femtoscopy is its potential to reveal signatures of the QCD critical point. Near the critical point, long-range correlations are expected, leading to a non-trivial, possibly non-monotonic energy dependence of the Lévy index $\alpha$ [16]. Motivated by this, we systematically analyze the energy dependence of femtoscopic source parameters within the EPOS4 event generator framework. This provides a crucial non-critical theoretical baseline for interpreting future experimental data and for exploring possible signs of critical behavior in particle correlations.
@@ 2.1 Distance distribution measurement and extraction of key parameters
In simulations, the pair distance distribution $D(\vec{\rho})$ can be directly obtained, where $\vec{\rho}$ represents the three-dimensional spatial separation vector of the particle pairs. And representative examples of the distributions are shown in Fig. 1, which are discussed in detail in Section 3.2. This allows the extraction of key source parameters such as the Lévy index $\alpha$, the source size $R$, and the correlation strength $\lambda$. In experimental measurements, however, the spatial distributions are not directly accessible. Instead, one measures the correlation function in momentum space (as a function of momentum difference $\vec{q}$), which encodes information about both the spatial distribution $D(\vec{\rho})$ and the pair wave function $\psi_{\vec{q}}(\vec{\rho})$:
@@
where $\vec{\zeta}$ is the integration variable, the superscript $T$ denotes the transpose, while $R^{2}=\mathrm{diag}\left(R_{\mathrm{out}}^{2},R_{\mathrm{side}}^{2},R_{\mathrm{long}}^{2}\right)$ denotes the diagonal matrix of Lévy-scale parameters defined in the Bertsch–Pratt “out-side-long” coordinate system [20, 21], neglecting cross-terms. Here the “long” direction is the direction of the beam, “out” is defined by the average pair momentum, and “side” is perpendicular to the previous two. The $2^{2/\alpha}$ factor appears due to $D(\vec{\rho})$ being defined as the autoconvolution of the Lévy-distributed single-particle source.
Figure 1: Pair distance distribution projections $D(\rho_{\mathrm{out}}^{\mathrm{LCMS}})$, $D(\rho_{\mathrm{side}}^{\mathrm{LCMS}})$, and $D(\rho_{\mathrm{long}}^{\mathrm{LCMS}})$ from EPOS4 simulation of 0-10% Au+Au collisions for the $k_{T}$ bin $0.225-0.275~\rm{GeV}/c$ at 7.7, 19.6 GeV, and 200 GeV collision energies.
In this study, for each energy, a total of 4000 minimum-bias events were generated using EPOS4 with its default X3FF Equation of State with cross-over transition and three flavour conservation [22, 23]. Out of these, the most central $0-10\%$ were selected via the predefined impact parameter limits given in EPOS4 (a multiplicity-based centrality selection was investigated as part of the systematic uncertainty estimates). The analysis was performed for identical pion pairs ($\pi^{+}\pi^{+}$ and $\pi^{-}\pi^{-}$). Tracks were selected with $0.15<p_{T}<1.0$ GeV/$c$ and $|\eta|<1$, following typical STAR kinematic settings. Unlike Ref. [24], where the angle-averaged distance distribution $D(\rho)=\frac{1}{4\pi}\int D(\vec{\rho})\,d\Omega$ was measured, but similarly to Refs. [25, 15, 26], we investigate the 3D pair distance distribution. To express the source coordinates in the Bertsch-Pratt coordinates of the longitudinally co-moving system (LCMS) [27, 28, 29], the components of $\vec{\rho}$ are computed from the particle positions in the laboratory frame as follows: [15]
@@
Here, the azimuthal angle $\varphi$ is defined by $\cos\varphi={K_{x}}/{k_{T}}$, $K_{0}$ and $(K_{x},K_{y},K_{z})$ denote the temporal and spatial components of the pair’s average four-momentum $\vec{K}$, $k_{T}=\sqrt{K_{x}^{2}+K_{y}^{2}}$ is the pair transverse momentum, and $(r_{x},r_{y},r_{z},t)$ are the particle coordinates in the laboratory frame. Note furthermore that in this paper we calculate $\vec{\rho}$ distributions in distinct bins of transverse mass $m_{T}=\sqrt{m^{2}+k_{T}^{2}}$, where $m$ is the mass of the investigated particle, in our case that of charged pions. Detailed derivations and conventions can be found in the Methods sections of Refs. [15, 25].
For each event, the distribution of relative distances $D(\vec{\rho})$ between pion pairs was constructed in the longitudinally co-moving system. The distance distributions were then fitted using the Lévy parameterization, with the fitting procedure following the methodology outlined in Ref. [15], to extract the parameters $\alpha$, $R_{\rm{out}}$, $R_{\rm{side}}$, $R_{\rm{long}}$, and $\lambda$. The fit quality was assessed through comparison of one-dimensional projections of theoretical fit curves and simulated distributions (similarly to Fig.5 of Ref. [15], and Fig.2 of Ref. [25]), ensuring that the main features of the source distributions are captured. The extracted parameters were also checked for consistency with the Lévy walk model and the expected core-core source behavior. The fitted distribution corresponds to the core-core source, consisting of pairs of pions produced close to the interaction point, i.e., primordial pions and the decay products of the shortest lived resonances  [25]. In addition, standard fit statistics $\chi^{2}$/ndf and confidence levels were evaluated to quantify the goodness of fit. Overall, the Lévy parameterization provides a satisfactory description of the freeze-out source across studied transverse mass bins and collision energies, supporting the interpretation of the extracted parameters.
@@ 2.2 Systematic uncertainties
The systematic uncertainties were estimated by varying several analysis conditions relative to the default configuration. The corresponding cuts applied for different sources are summarized for each setting below.
(1) Number of averaged events. Event-by-event studies show that Lévy-shaped sources already appear in individual EPOS events [24, 30], indicating that the non-Gaussian features are not caused by event averaging. However, due to limited pair statistics in single events, events were merged in the present analysis to obtain stable Lévy parameter extraction. It was carefully investigated and confirmed that the extracted parameters converge at a sufficiently large event number. For different collision energies, the number of merged events ($N_{\rm events}$) was varied as follows (the middle value is taken as default):
  • 1.
@@
  • 3.
    130 and 200 GeV: (70, 100, 130)
(2) Pair momentum difference. In quantum-statistical correlations, the signal appears at limited values of relative momentum $Q_{\mathrm{LCMS}}$. Based on the observed $\sqrt{m_{T}}$ scaling of femtoscopic correlation widths  [27, 31], we utilize pairs up to a maximal momentum difference of $Q_{\mathrm{LCMS}}^{\mathrm{max}}=c\cdot\sqrt{m_{T}\cdot 0.001\text{ GeV}}$ where the coefficient $c$ was varied for different collision energies (again the middle value is the default):
  • 1.
@@
(3) Fit range in $\rho$. The stability of the extracted source parameters was examined by varying the fitting range in $\rho$:
  • 1.
@@
(4) Centrality definition. Centrality was defined using either the impact parameter (bimbim) from the EPOS4 definition or the charged-particle multiplicity ($N_{\mathrm{ch}}$). The two-dimensional ($bim,N_{\mathrm{ch}}$) distributions of the dataset were analyzed to define centrality bins according to the accumulated event fractions (0-10%, 10-20%, etc.). The centrality boundaries extracted from the impact parameter distribution are further found to be consistent with those provided directly by EPOS4. Alternatively, we also considered a multiplicity-based event selection. Overall, the choice of centrality definition contributes less than 2% to the systematic uncertainty of all extracted femtoscopic quantities.
A systematic uncertainty study considering the above points was performed for all collision energies and each $m_{T}$ bin. The results show that, in the region $m_{T}<0.7~\mathrm{GeV}/c^{2}$, the systematic uncertainties of the parameters exhibit no significant dependence on collision energy or $m_{T}$. In this region, the uncertainties fluctuate mildly around a similar level, with typical magnitudes of about 0.5%–3% when averaged over $m_{T}$ bins. For $m_{T}>0.7~\mathrm{GeV}/c^{2}$, the uncertainties increase noticeably, reaching up to 13%, primarily due to the reduced pair statistics at higher transverse momentum. When averaged over all $m_{T}$ bins, the total systematic uncertainties decrease with increasing collision energy. Consequently, the physical interpretation of the $m_{T}$ dependence in this work focuses primarily on the region $m_{T}<0.7~\mathrm{GeV}/c^{2}$, where the results are more robust and the uncertainties are better controlled.
@@
Table 1: Relative systematic uncertainties of Lévy parameters. (a) Relative systematic uncertainties from each source at selected energies: 7.7, 19.6, 200 GeV. (b) Summary of total relative systematic uncertainties over all analyzed energies.
@@
$\sqrt{s_{{}_{NN}}}$ Source $\alpha$ $\lambda$ $R_{\mathrm{out}}$ $R_{\mathrm{side}}$ $R_{\mathrm{long}}$
7.7 $N_{\rm events}$ 0.6% 0.3% 0.2% 0.3% 0.5%
GeV $Q_{\mathrm{LCMS}}^{\rm max}$ 0.7% 0.1% 0.2% 1.0% 2.2%
$\rho_{\rm fit}^{\rm min/max}$ 0.4% 0.2% 1.0% 0.1% 0.7%
Centrality 0.9% 0.2% 0.1% 0.6% 0.8%
Total 1.3% 0.4% 1.1% 1.2% 2.4%
19.6 $N_{\rm events}$ 0.3% 0.1% 0.01% 0.1% 0.2%
GeV $Q_{\mathrm{LCMS}}^{\rm max}$ 0.5% 0.03% 0.4% 1.0% 1.7%
$\rho_{\rm fit}^{\rm min/max}$ 1.3% 0.34% 0.4% 0.1% 0.8%
Centrality 0.5% 0.2% 0.4% 0.5% 0.2%
Total 1.5% 0.5% 0.7% 1.1% 2.0%
200.0 $N_{\rm events}$ 0.1% 0.1% 0.1% 0.1% 0.1%
GeV $Q_{\mathrm{LCMS}}^{\rm max}$ 0.4% 0.4% 0.2% 1.0% 2.0%
$\rho_{\rm fit}^{\rm min/max}$ 0.5% 0.1% 1.0% 0.3% 0.9%
Centrality 0.2% 0.2% 0.6% 0.4% 0.3%
Total 0.7% 0.5% 1.1% 1.1% 2.2%
(b) Summary of the average total relative systematic uncertainties (in %) of the Lévy parameters over all studied $m_{T}$ intervals ($m_{T}\approx 0.245\text{--}0.663~\mathrm{GeV}/c^{2}$), for the analyzed $\sqrt{s_{NN}}=7.7\text{--}200$  GeV energy range.
$\sqrt{s_{NN}}$ (GeV) 7.7 9.2 11.5 14.5 19.6 27 39 62.4 130 200
$\alpha$ 1.1 1.3 1.2 1.1 1.5 1.5 1.2 1.24 1.3 1.0
$\lambda$ 0.6 0.6 0.5 0.6 0.7 0.5 0.6 0.6 0.5 0.5
$R_{\mathrm{out}}$ 1.2 1.0 0.9 1.1 1.0 0.9 0.9 1.0 0.9 1.1
$R_{\mathrm{side}}$ 1.2 1.1 1.1 1.2 1.0 0.9 1.0 1.0 1.0 1.0
$R_{\mathrm{long}}$ 2.6 2.6 2.4 2.3 2.1 2.2 2.0 2.2 2.2 2.1
@@ 3 Results and discussions
In this section we present the results for the femtoscopic source parameters as functions of collision energy, focusing on three main aspects: source shape ($\alpha$), source size ($R_{\text{out}},R_{\text{side}},R_{\text{long}}$), and correlation strength ($\lambda$).